by TEI TEI

03/15/07 – TTC & Bill Scholl setup a meeting to check VFD operation for next Wednesday Morning.

02/19/07 – Project is complete. Contractors have finished all the punchlist items and the design changes have been fixed. TTC & Bill Scholl will balance the chiller as soon as the cooling towers are filled up and ready to run in 2007. O&M have been received and completed and sent to Tony Matre.

01/24/07 – Triton may start equalizer line on cooling towers. County may start to drain the cooling towers this Friday for this work. TTC and Bill discussed QCE providing power to the VFD but it being manually dialed into a HZ that equals the correct pressure drop across the chiller at full flow. This will be done after Trane confirms the dP and procedure that TTC wants to use to balance the system.

01/17/07 – Tony requested the project be closed out and archived. TEI sent the contractors this imperative. TEI will work to get paperwork completed in one week.

12/26/06 – Bill Scholl reported that the chiller is repaired but now the RM is going off into 3rd stage alarm. We decided just to turn off the RM until A&M can get to the sire to recalibrate the unit and verify it is operational again. This will be done the first week of January along with the new programming of the Chiller plant per last months instructions. The VFD is in town and QCE will be scheduling the install the first week of Jan 2007. QCE may be out earlier to mount the unit and wire it up for the factory rep to program later. BS and TTC discussed and are in agreement with this procedure. Once the VFD and pump are up and running we can manually dial-in the pump to 800 gpm (or a factory agreed PD across the chiller barrel) and then set it and forget it. That should wrap up that project.

12/22/06 – QCE reports that the VFD has arrived. We are scheduling installation immediately. TTC contacted Triton about pump vibration and resolution and will pass along information to County.

12/18/06 – VFD has been ordered. When it comes in, QCE will install. The chiller had a R leak that Trane is fixing. The oil leak to the bearing was found not to be an issued per the local Trane office so the bearing will not need to be replaced.

12/01/06 – Monthly Update. Triton removed the motor from CHWP-1 and sent it out for rebuild. According to Grady, it still vibrates but a lot less. Triton is going to have Weber-Huff come over to see the motor run independent of the system (un-coupled) and prove the motor has a balance issue. We are hoping they either repair or replace the motor. The motor fix is independent of the pumping issue where we have decided to put a VFD on the CHWP-1. AKM has secured a PO for the replacement and TTC has talked to Greg Chambers about ordering the GE VFD and he has. It is a stock item that should only take a week or two to come in. We will then schedule the VFD for installation. Once the VFD is installed, we will balance the constant volume system for 800 gpm through CHLR-1 and mark that position on the chiller, this should be the standard run % of the pump at all times. The remaining allowance usages have been coordinated with Grady at Triton as such: Triton will put the equalizer in between the towers. They will used companion flanges per TEI request and use 6″ SCH 80 PVC piping. The will use minimal elbows on the system to cut down on static but will drop the EQ line to the floor and back up to the basin on both towers. The remaining balance will be paid directly to A&M to have the backup CHLR-2 automated as the backup chiller and amend the economizer sequence such that the chiller plant will stay off line while the air handlers are in economizer mode (currently the chiller is only started by OA temp). The County will move the separator and have A&M pipe fitters provide the taps for the separator. Once all this work is done, the project should be complete. TTC has submitted payapps for the contractors but TEI has to write a change order to QCE for the VFD and allowance usages to Triton for allowances.

10/31/06 – Triton came to the site to check the alignment of the pump sled. By the end of the day they removed the motor for warranty repair. I do not know how long it will be gone but will keep everyone apprised of the dates once they are known. TTC has contacted Triton about adding the flow orifice or the additional balance valve to the job as well. Triton provided a price to do the equalizing line work which is more than expected. TTC passed it on to Tony.

0916/06 – TTC visited the 237 site again on 09/11 to review pump status conditions. I measured the actual ft of hd of the whole system and found it to be 65. TEI sent email messages to Tony, Bill and Dave advising of the existing conditions. There are a couple of things we can do to make the pump better match the existing conditions of the building (e.g trim the impeller, add a VFD, buy a different motor and pump. Weber-Huff reported that just reducing the HP is not a viable option because of the volute on our pump is too big for this small of head in the new system). Tony asked for meeting to discuss the situation and TTC will set that up. Final inspection is complete but the inspector wants to see the refrigerant evacuation system work. TTC advised Triton to coordinate with A&M and get that done for the County. TEI has already seen the system work a couple of times.

09/01/06 – After visiting the site with the Hydrodata meter, TTC contacted the vendors and explained in greater detail the operation of the new pump and chiller. The pump appears to have more head than is needed for this system. At 86 ft hd we have 925 gpm of flow. The design conditions are 800 gpm at 110 ft of hd. TTC offered the County several options on managing the issue. TTC emailed Bill S and Tony M with the recommendations.

07/14/06 – Bill called about training and TTC directed contractors to provide this training and coordinate it with Bill Scholl. Then provide closeout docs to finalize the project. TTC asked for direction from County on using allowance money for project before closeout. TEI would like to close this project and give back extra SOSS money to Dave Woodside before it is used up.

06/27/06 – Tony reported the Bill reported that the new drainage system is working and now water leaks are currently flooding the floor at the building. This good news. We can now close out the project as all other items have also been resolved.

06/26/06 – TTC stopped at the 237 building and put in a new drainage system on the cooling towers consisting of 1″ piping and funnels to catch the weep hole drainage of the cooling towers.

06/06/06 – TTC was on site today to conduct the refrigerant monitor system test. A&M ran the test early and Bill Scholl thought it worked. No one wanted to take a chance of setting off the horns at 7am so we let the early test be the test. TTC witnessed that both sets of pumps backed up the other pumps in about 30 seconds each. The 2nd chiller is truly a backup manual chiller that is enabled and disabled by the front end. The system while rotating chillers runs an extra 10 minutes to remove residual heat from the equipment before completely shutting down. TTC set the triple duty valve to 30% open and brought the CHWP-1 noise to down to a very good level. The pump appears to be running fine with no cavitation sounds and the air handlers appear to be getting enough flow and temperature to work properly. Later during the morning, the RM went into alarm at the second level with strobes on but the front end did not report this 2nd level alarm. Bill Scholl said that this was similar to a previous test where the false readings were from the method the contractor used to test the system. A&M stayed until they fixed the problem with a wiring change and talking to the factory. TTC removed the pans from CT-2 and they were water logged and water was all over the floor anyways. We added some plastic gutter drains for a TEMPORARY TEST to channel the water from the fans to the floor drains. We are not complete and TTC set the fire alarm off in the building by running the shop vac to remove excess water off the floor (that’s a pretty touchy smoke detector). We have to add more plastic drainage to channel the water even more directly to the floor drains. We are trying to dry out the penthouse to pinpoint exactly where CT-2 is leaking from. NOTE: even though the towers have small weep hole leaks like they are DESIGNED TO have by Baltimore Air Coil, my promise that they won’t flood the room from overflowing is STILL HOLDING TRUE.

05/22/06 – TTC visited the site and performed the controls punchlist for the project. Only minor items were found at this punchlist. TEI needs to coordinate a field trip for testing the backup systems on the chillers, pumps, refrigerant monitor system in live action. TEI to coordinate electrical punchlist on this job as well. The equipment is up and running and Bill Scholl seems to have a good handle on how to run the new system as yesterday when it was only 64F outside the new chiller was disabled and the smaller chiller was in service. A&M provided a control point to handle this control.

05/01/06 – TEI has been working to get the Controls finished on the job. A&M wrapped up work last week. Trane and its XIF file is not communicating properly with the front end and once that is complete the job could be finished. CHWP-1 is vibrating too much and Rick Huff believes it must be the pressure drop across the pump as it is readying way too low. TTC will check this out.

04/17/06 – TEI confirmed balancing with contractor on 04/18/06 at the building. Triton, Approved Products and TEI will be on site.

04/11/06 – TTC sent an email out to contractor to inquire about punchlist day, balancing day and controls for project. TTC instructed A&M to modify their existing sequence to match the intent of the specs for the pumping arrangement as the pumps are lead/lag when in reality they are supposed to be matched to the chillers they serve. Pump vibration is still excessive and TTC would like that looked at as well. Tony asked for a control meeting on the cooling towers and TTC gave him Thursday/Friday dates for this meeting.

04/05/06 – TTC talked to A&M about project progress and they reported that the controls can realistically be expected to be complete by April 14, 2006. The electricians and programmers are on site right now completing the work. A&M asked about how to program the refrigerant exhaust fan before or after the local hand/off/auto switch, TEI stated that the RM must be able to turn the exhaust fan on even if the HOA switch is in the “off” position. A placard will be made to designate that if total disconnect of the EF-1 is required, the breaker for the fan must be shutdown and locked out. It should be noted that this will bypass the automatic controls to evacuate the refrigerant from the space in case of leaks but keeps the workers completely safe otherwise. The chiller has been tested under low load conditions. Trane would like to make an secondary visit to the site to test it under full load conditions as the weather permits. TEI will schedule the punchlist work for mechanical and electrical within the next 5 days. TTC will punchlist controls after they are complete. As far as TEI knows at this point, either chiller can be used to cool the building but CHLR-2 is presently being used until Trane makes it return visit. Training is to be coordinated between the County and Vendors as needed at owners discretion.

04/04/06 – Bill Scholl called today to talk about control sequences on new cooling towers. TTC provided the history of the cooling tower plant that he was aware of. Here is the recap: The old cooling tower design allowed for one bank of a cooling tower to run independent of the other, this allowed moisture laden water to re-infiltrate the “off” tower and condensate in the fan housings and leak onto the penthouse floor and then onto the ceiling of the floor below. TEI changed the existing controls to force both banks of a tower to come on together to prevent this from happening (this may have changed when Fosdick modified the control sequence in the AHU job – Bill reported that the cooling towers were no longer coming on as one unit). Our new control sequence calls for actuators to be installed to control the backdraft dampers on the towers to open and close when the towers are in operation. As these are really four towers operating as two, let me use the declarations of CT-1A and CT-1B to represent one cooling tower and CT-2A and CT-2B to represent the second cooling tower. In the new design and sequence, TEI is asking the cooling towers to be put into operation when the chiller is enabled and either one of the condenser water pumps are enabled. Based on the outside air temperature, a reset temperature for the condenser water system will be deemed as a setpoint. As the condenser water temperature rises CT-1A fans will come on and open its cooling tower dampers along with the intake dampers to bring the temp back down to setpoint. If these fans can’t keep up CT-1B fans will come on to keep up and its dampers will open. The operator will select which towers to use in which order as 1,2,3,4 or 2,3,4,1 or 3,4,1,2 or 4,1,2,3 where 1 = 1A, 2=1B, 3=2A, 4=2B. There is no equalizer line between the sets of towers because the size of an equalizer line to let enough water flow between the basins without residual static was either 12 or 14 inches round – not practical PLUS you have a constant volume system, which once balanced shouldn’t have flow problems. The new strainers should keep the tower cells clean and the flow equal between them. Also when checking the design of the towers – they are sized for 600 tons on chillers – we are only using 400 tons – there should not be a problem with condenser water volume. The overflows are now piped directly to the floor drains so if for some reason, the basins overflow, the water should go down the drain, not onto the floor. The backdraft dampers put on the towers at the end of the previous public works project were all that could be afforded at the time of purchase. The actuator system that A&M is adding to the job at this time should automate these dampers one tower at a time (we don’t specifically tell A&M how to automate them, only that they need to be automated, so we will review their design when complete – or they can ask questions if needed). We just want to see them work when complete. If all else fails, you can always block the dampers open like you have been to get full performance if that becomes and issue. Under normal operation, we don’t envision any problems with the cooling tower operation under constant flow. If someone walked up the outlet of the towers and slammed the valves shut and all the CT basin’s overfilled before the pump suction cavitated, maybe there could be an issue – but we don’t see that happening on accident.

03/30/06 – Triton will have a tech at the building to check pump alignment on the new pumps on Friday. The backup chiller system was operational today and seemed to be working fine. The new chiller needs to be load tested under warmer days but was fine when started in the preliminary stages. TEI still wants to see the cooling towers balanced under full flow through the new chiller and pumps. QCE had to come in and change the phases on the chiller before it was started up for testing.

03/29/06 – TTC stopped at building today. Backup pumps and chiller are ready to start. County found a enable switch on an old control panel that turned the second pumps on. A&M is installing the RM system. Jerry Smith was working on that today and reported that the new actuators would soon be wired up on the new chiller. The Trane rep had the modules for the DPS switches in hand was hooking them up. He stated that he may be able to start the chiller today and begin to run it through its initial paces. The temperatures are supposed to be higher on Thurs/Fri so we may need the backup chiller. Mike Iverson from A&M is supposed to stop at the building this morning as well.

03/28/06 – Bill reported today that PSD switches are not installed. TTC sent out Const Mtg notes #4 to team. Chiller might be started this Thursday. TTC checking on status of controls with A&M services. LON cable is pulled, A&M is installing the cooling tower dampers now and they are waiting on some XIF files from Trane for LON integration into the new chiller.

03/24/06 – Bill called today to report that there was a problem with the oil filter on the chiller not sealing. Apparently it was damaged during installation by being stepped on. Later, Trane was able to put the replacement filter on and prove it sealed perfectly. We are on schedule to start the chiller next Tuesday or Wednesday – we have to wait and see when Trane finishes to be sure. TTC will call Tony to let him know when we know. A&M reported today that they should be onsite next week to finish installation and programming.

03/23/06 – Grady reports that chiller should be 100% assembled by end of business today with any luck. TTC asked A&M for controls schedule for completion.

03/21/06 – The County reported today that the Chiller piping is complete and we are filling the system. Triton is waiting on a pressure relief valve for new chiller and then we will fill it. Queen City Electric has the wiring pulled to chiller as of today. They will be back tomorrow to terminate wiring. A&M have been pulling control wire. I am hoping to get power on chiller tomorrow and then call Trane back in to set up chiller. TTC is going to stop at the building on Wednesday to check over the progress. TTC is checking on chiller startup date with Trane and Triton. The chiller will not start up on Wednesday because the DP switches aren’t here yet. Soon.

03/20/06 – TTC emailed A&M back with alarm settings for the Sentech monitor. First phase = LCD on Sentech, Second Phase = Strobes and LCD, Third Phase = Horn, Strobes and LCD.

03/15/06 – TEI spent hours today trying to figure out if the chiller really needs external pressure switches? Once the vendor replies to these questions, TEI will be able to give Triton direction on how to handle this question. Trane emailed me back and either Triton or Trane will provide the PSD on the project, TTC is letting Triton decide.

03/08/06 – Bill called today to inquire about insulation on the evap barrel on the chiller. TEI reported that it should be added after the chiller is put back together. TTC distributed the Const Mtg 03 notes.

02/27/06 – Bill Scholl left a message on my phone today stating that Tony said Facilities will move the light in the penthouse and paint the housekeeping pad at the 237 building as part of this job and Triton will not have to do it. Grady reported that the chiller arrived right on time and has begun its disassembly per the project schedule. We have a meeting this Friday for thsi job.

02/23/06 – Bill Scholl called today to inquire about paint color on the chiller pad pursuant to our last construction meeting. TTC told bill to tell Pat from Triton that the color should be gray and match the floor sealant color the County used when the mechanical room was upgraded. The actual painting of this pad was not in the scope of the project but can be covered out of the allowance money on this project. TTC visited the site earlier in the week and reviewed new valve install with Pat from Triton, we coordinated the on sight drawing for what was actually being installed in the field. Pat and Terry discussed the separator install and Triton agreed to pipe the separator in and out of the condenser water supply pipe and maintain at least six feet separation between inlet and outlet of separator. Triton also pointed out that one of the make-up water valves was damaged and actually “locked” on. There is no telling how long it has been like this but may explain the unusual high water usage at the building last year. TTC asked Triton to provide a cost to replace all 4 valves before we button the towers back up for service. TTC also asked Pat to clean up the drains going into the floor drain so that any overflow from the towers goes inside the existing FD’s and not out onto the floor. Triton was well on their way to having both strainers installed in the header feeding the cooling towers. TTC had to instruct Triton to install spool pieces on each side of the strainers because the strainer interfered with flapper operation on the valves. There is an existing light fixture that is way to close to the strainers and should probably be moved (does the County trades want to move or should I contact QCE about relocated the penthouse light?). Control submittals were received and marked up by TEI and mailed back to Grady Reed today.

02/22/06 – TTC emailed the team to say that A&M’s lockout on the control valves is acceptable. No additional valves are need for this job.

02/16/06 – TTC sent out request for additional valves on chillers per HVAC Trades request. TTC sent out request for decision on separator rework per HVAC trades request and reminded everyone about Friday’s construction meeting.

02/06/06 – TTC rescheduled the next const mtg for for 02/17/06 at 1:30 p.m. ThermalTech received submittals for the strainers, insulation, valves, gages and meters, refrigerant monitor. TTC will work on these this week.

01/23/06 – TTC distributed Construction Meeting # 01 Minutes.

01/18/06 – TTC found out today that the Trane has finalized the ship date of 02/24/06. This basically gives Triton about 2 weeks to install the chiller and meet the 03/15/06 deadline. TTC questioned everyone and is trying to get written buy-in that the project schedule can be met.

01/11/06 – Pre Construction meeting notes distributed. Triton is reporting that chiller won’t be shipped until 03/10/06. TTC has already placed emails and phone calls to see why the date is so late in the year. This is much later than we would care to have for this project and although a backup chiller will be ready to handle the load, we wanted to punchlist this job on 03/01.

01/06/06 – Pre Construction meeting was held. TTC will type up and send out meeting notes soon. The county asked for employee list and security checks. Triton may be on site next Wednesday for field review and to get familiar with building. Contractors want 8 additional sets of dwgs for this project. TE will have sent to contractors.

12/19/05 – TE finished the chiller submittals and sent them to Triton for ordering. After contracts are approved and PO are received, a pre const mtg will be held.

10/14/05 – TE delivered the final contracts to Kathy Bushbacher for the project. The prosecutor marked up and returned and TE reprinted contracts and delivered additional drawings to JFS so that this project could stay fast tracked. TE delivered 8 new contractors within 1-1/2 hours of the phone call and TTC hand delivered to JFS.

10/13/05 – The bids opened on 10/0705 and after much internal work inside ThermalTech verifying the LCC equations, TE recommended the Trane Chiller and Triton Services as the installing contractor. Queen City Electrical was the electrical low bidder on this project. TE is writing the contracts for delivery to Dave Woodside today so that he can turn them in to the HamCo Prosecutor for review.

09/29/05 – Two critical equations are missing from the life cycle cost equation sheet and TE must extend the bid form to allow contractors to enter this data. The bid is being extended one week by Addendum # 02.

09/27/05 – York called today to ask about chiller retrofit on CHLR-2. We discussed how the County will or won’t accept the alternate to change out the refrigerant.

09/26/05 – Addendum 01 has been issued for this project. Bid Opening is still scheduled for this Friday.

09/06/05 – Project was delivered for bidding to Purchasing and Ohio Blue. Prebid is set for 09/16 at 1:30 p.m.

08/31/05 – TTC sent out an email over viewing all the construction and production changes to the final design at the 237 building. The County is seriously concerned about the cooling tower overflow issue and TTC assured Dave Woodside that the construction project will address this issue through cleaning, balancing and repiping the cooling towers. The new system is constant volume and once balanced there should be no more remaining balancing issued between the separate cells on the towers. TE will have the mechanical contractor take care of all these issue during the construction of the cooling tower project.

08/01/05 – TE spent the last two weeks doing field work and finalizing the design on this project. Construction documents should be ready to review at mid August and be ready for bidding. We are going with a 45 degree chilled water temperature and a 12 degree delta T to meet existing conditions at the site and the new ahu coils in use. TE will schedule the final review within the week and go over the final plan for the project and get consensus on all that we are doing.

06/14/05 – TE kicked off a design meeting at the site. We reviewed the scope, bidding process and alternates to the project. See the meeting notes below.

06/13/05 – TE set to meet client for kickoff meeting at site.